OpenStudy (blasofia98):

Two similar hexagons have areas of 36 square inches and 64 square inches. The ratio of the corresponding sides of the hexagons is _____.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Look at the figure above. There are 2 squares. One is 1 by 1. The other in is 3 by 3. The ratio of the lengths of the sides of the larger one to the smaller one is 3 to 1. Ok so far?

OpenStudy (blasofia98):

yes

OpenStudy (mathstudent55):

Now can you find the areas of the two squares?

OpenStudy (mathstudent55):

Area of square = square of length of side Side length = 1. Area = 1^2 = 1 Side length 3. Area = 3^2 = 9 The areas are 1 and 9. Ok?

OpenStudy (mathstudent55):